OpenStudy (anonymous):

Which of the following equations is of a parabola with a vertex at (0, -3)? y = (x - 3)2 y = (x + 3)2 y = x2- 3 y = x2+ 3

Parth (parthkohli):

Something that'd look like this \(\Rightarrow y = (x - 0)^2 + 3 \)

Parth (parthkohli):

x - 0 = x \(\Rightarrow y = x^2 + 3\)